### v1

# The Feasibility of a Spacecraft Flyby with the Third Interstellar Object 3I/ATLAS from Earth or Mars

Abstract
We investigate the feasibility of a spacecraft mission to conduct a flyby of 3I/ATLAS, the third macroscopic interstellar object discovered on July 1 2025, as it traverses the Solar System. There are both ready-to-launch spacecraft currently in storage on Earth, such as Janus, and spacecraft nearing the end of their missions at Mars. We calculate minimum $\Delta V$ single-impulse direct transfer trajectories to 3I/ATLAS both from Earth and from Mars. We consider launch dates spanning January 2025 through March 2026 to explore obtainable and hypothetical mission scenarios. Post-discovery Earth departures require a challenging $\Delta V\gtrsim 24$ km s-1 to fly by 3I/ATLAS. By contrast, Mars departures from July 2025 - September 2025 require $\Delta V\sim 5$ km s-1 to achieve an early October flyby — which is more feasible with existing propulsion capabilities. We discuss how existing spacecraft could be used to observe 3I/ATLAS and how spacecraft at other locations in the Solar System could be repurposed to visit future interstellar objects on short notice.
[…]
show]yaginuma@msu.edu
1 Introduction
The discovery of 3I/ATLAS (L. Denneau et al., 2025) increases the interstellar object sample by $50\%$, building on the previous discoveries of 1I/‘Oumuamua in 2017 (G. V. Williams et al., 2017) and 2I/Borisov in 2019 (G. Borisov et al., 2019). On 2025 July 1, the Asteroid Terrestrial-impact Last Alert System (ATLAS) survey (J. L. Tonry et al., 2018) discovered the small body A11pl3Z  [^1] https://www.minorplanetcenter.net/mpec/K25/K25N12.html. Fits to the ATLAS astrometry revealed that this object has a perihelion distance $q\sim 1.36$ au, orbital eccentricity of $e\sim 6.2$, inclination of $i\sim 175^{\circ}$, and a hyperbolic velocity of $v_{\infty}\sim 58$ km s-1, clearly indicating an interstellar origin. This object is only the third confirmed macroscopic interstellar object and therefore presents an unprecedented opportunity to characterize a planetesimal from outside of the Solar System.
Preliminary observations of 3I/ATLAS revealed a weakly active nucleus (D. Jewitt & J. Luu, 2025; M. R. Alarcon et al., 2025; C. O. Chandler et al., 2025) with a reddened reflectance spectrum and an absolute magnitude of $H_{V}\sim 12$ (D. Z. Seligman et al., 2025; B. T. Bolin et al., 2025; C. Opitom et al., 2025) although the nucleus size is uncertain (A. Loeb, 2025). Its visible and near-infrared reflectance is red and featureless, matching the spectral slopes of D-type asteroid and active Centaurus (M. Belyakov et al., 2025; T. Kareta et al., 2025; R. de la Fuente Marcos et al., 2025). 3I/ATLAS’s high excess velocity and Galactic velocity radiant imply an old kinematic age of 3–11 Gyr (A. G. Taylor & D. Z. Seligman, 2025), consistent with an origin from the thick disk of the Milky Way (M. J. Hopkins et al., 2025; S. Kakharov & A. Loeb, 2024).
The first two interstellar objects exhibited remarkably different characteristics. 1I/‘Oumuamua displayed no cometary activity (K. J. Meech et al., 2017; Q.-Z. Ye et al., 2017; D. Jewitt et al., 2017; D. E. Trilling et al., 2018) but its trajectory exhibited a significant nongravitational acceleration (M. Micheli et al., 2018). In stark contrast, 2I/Borisov displayed clear signatures of active coma (D. Jewitt & J. Luu, 2019; A. Fitzsimmons et al., 2019; Q. Ye et al., 2020; A. J. McKay et al., 2020; P. Guzik et al., 2020; M.-T. Hui et al., 2020; Y. Kim et al., 2020; G. Cremonese et al., 2020; B. Yang et al., 2021). 1I/‘Oumuamua and 2I/Borisov had lower excess velocity than 3I/ATLAS, corresponding to younger kinematic ages albeit with large uncertainty (E. Mamajek, 2017; E. Gaidos et al., 2017; F. Feng & H. R. A. Jones, 2018; F. Almeida-Fernandes & H. J. Rocha-Pinto, 2018; T. Hallatt & P. Wiegert, 2020; C.-H. Hsieh et al., 2021).
The discovery of these interstellar objects has ignited significant interest in sending dedicated spacecraft to measure their properties in situ (e.g., D. Seligman & G. Laughlin 2018; K. Moore et al. 2021). Space-based measurements of interstellar object physical and chemical attributes — such as isotopic signatures and volatile compositions — would profoundly enhance our understanding of extrasolar systems at both broad and narrow scales. Compositional indicators such as the H2O content — when combined with the kinematics and galactic history of a given object — can provide insight into planet and star formation rates throughout the galaxy (C. Lintott et al., 2022; M. J. Hopkins et al., 2023, 2025). On the other hand, the carbon-to-oxygen ratio of an interstellar object approximately traces its formation location in a protoplanetary disk (D. Z. Seligman et al., 2022b). Since analogous objects may be responsible for enriching short-period exoplanet atmospheres (D. Z. Seligman et al., 2022a), such data could provide insight into the delivery of volatiles and organics to exoplanets.
Many spacecraft have conducted close-range remote observations of comets and asteroids in the Solar System, such as the Deep Impact (M. F. A’Hearn et al., 2005), Giotto (R. Reinhard, 1986, 1987), NEAR Shoemaker (L. Prockter et al., 2002), Rosetta (K.-H. Glassmeier et al., 2007), DART (A. F. Cheng et al., 2023), and Hera (P. Michel et al., 2022) missions. Moreover, a number of sample-return missions have delivered extraterrestrial material back to Earth, e.g., Stardust (D. Brownlee, 2014; D. E. Brownlee et al., 1996), Hayabusa (M. Yoshikawa et al., 2021), Hayabusa2 (S.-i. Watanabe et al., 2017), and OSIRIS-REx (D. S. Lauretta et al., 2017).
Recently, Comet Interceptor — an ESA F-class mission — has been approved and is currently in its implementation phase (C. Snodgrass & G. H. Jones, 2019; G. H. Jones et al., 2024). Designed as a flexible mission architecture, Comet Interceptor will be stationed at the Sun–Earth L2 point, awaiting the discovery of a dynamically new comet or other pristine small body to intercept. Unlike previous missions, Comet Interceptor will employ a multi-spacecraft configuration to enable simultaneous multi-point observations of the target. This mission’s primary objective is to investigate a long-period comet entering the inner Solar System for the first time. However, if early warning conditions permit, it could be redirected to perform the first in situ exploration of an interstellar object.
The discovery of 1I/‘Oumuamua and 2I/Borisov rapidly spurred development on an interstellar object mission concept. Initial studies (e.g., D. Seligman & G. Laughlin 2018) emphasized the urgency of preparing interception missions with minimal latency but determined that it was feasible to detect and reach future interstellar objects with conventional chemical propulsion. Further studies such as Project Lyra examined the feasibility of direct missions to 1I/‘Oumuamua using near-term technologies, including solar Oberth maneuvers and high-energy launch architectures (A. M. Hein et al., 2019; A. Hibberd et al., 2021; A. Hibberd, 2023). These analyses demonstrated that with optimized trajectories and sufficient launch energies, flyby or rendezvous missions could be viable even years after 1I/‘Oumuamua’s perihelion. Alternative mission design proposals include solar sail architectures capable of prolonged loitering and rapid-response interception of future interstellar objects (A. M. Hein et al., 2019; D. Garber et al., 2022; D. Miller et al., 2022), or a distributed statite cluster for high-$\Delta V$ response scenarios (D. J. Hoover et al., 2022). Evaluations of optical and autonomous navigation performance during high-velocity flybys emphasize the importance of on-board image processing and neural network guidance to maintain targeting accuracy at encounter speeds exceeding 60 km s-1 (D. Mages et al., 2022; B. P. S. Donitz et al., 2023). A complementary analysis focused on the tradeoffs between target detectability, spacecraft storage strategies, and imaging resolution advocated for rendezvous missions to 1I/‘Oumuamua-sized targets (A. Siraj et al., 2023). The “Bridge” New Frontiers-class flyby mission concept would intercept an interstellar object with a dedicated science payload including a guided impactor, ultraviolet/visible/infrared spectrometers, and a high-resolution camera (K. Moore et al., 2021). Additional studies explored broader science cases and mission architectures for interstellar object investigation, including rapid-response trajectories and programmatic feasibility (B. P. Donitz et al., 2021), the astrobiological relevance of interstellar objects as potential carriers of organic material and the propulsion technologies that could enable their exploration (M. Lingam et al., 2023), and the practical constraints and feasibility of near-term rendezvous missions to interstellar objects using advanced but achievable propulsion architectures (D. Landau et al., 2023).
Here, we investigate possible missions that could reasonably intercept 3I/ATLAS in the months following its July 2025 discovery. We specifically explore utilizing existing missions near Earth or Mars and consider favorable minimum distances to 3I/ATLAS. This paper is structured as follows — in Section 2, we detail an algorithm designed to determine direct transfer trajectories and launch dates. In Section 3, we calculate minimum-energy mission trajectories to 3I/ATLAS from Earth achievable in the months following its discovery. In Section 4, we identify additional key launch dates and trajectories for a flyby mission to 3I/ATLAS from Mars. Finally, we conclude in Section 5 with a summary of our results and a final discussion of our suggestions for immediate implementation of an intercept mission with 3I/ATLAS.
2 Identification of Optimal Trajectories
In this section, we describe the algorithm that we implemented to find minimum-energy trajectories to 3I/ATLAS. This algorithm is based on a methodology to calculate the minimum-energy trajectory between two elliptical orbits (H. Leeghim, 2013), which was later expanded to include hyperbolic trajectories in Section 4 of D. Seligman & G. Laughlin (2018).
The algorithm first solves Kepler’s equations as a function of time given a departure date $\tau_{0}$ that is sampled over a 400-day interval. For the remainder of this paper, a subscript of 0 denotes departure and I/T denote interceptor/target. The flight time is a free parameter in the optimization problem and is denoted by $\Delta\tau$. We obtain the target’s heliocentric position from JPL Horizons, and then simplify the boundary-value problem by enforcing only the interceptor’s universal variable residual. For each candidate departure and flight duration, $\Delta\tau$, we extract the interceptor’s heliocentric state ($\vec{r_{0,I}}$, $\vec{v_{0,I}}$) and compute the target’s position $\vec{r_{T}}$ at arrival. For every pair of both (i) departure date and (ii) arrival date, we solve the minimum-energy single-impulse trajectory by applying the two-body Lambert problem using the universal variable method. With this method, the function $\eta$, which indicates how well the solution fits the Kepler’s equations of motion, may be written as:
$\begin{split}\eta\big{(}\chi,\tau,\vec{r}_{0},\vec{v}_{0}\big{)}=\lVert\vec{r}_{0}\rVert\chi-\sqrt{GM_{\odot}}\Delta\tau\\ +\frac{\vec{r}_{0}\cdot\vec{v}_{0}}{\sqrt{GM_{\odot}}}\,\chi^{2}\,C\bigg{(}\frac{\chi^{2}}{a}\bigg{)}\\ +\left(1-\frac{\lVert\vec{r}_{0}\rVert}{a}\right)\chi^{3}S\bigg{(}\frac{\chi^{2}}{a}\bigg{)}\,.\end{split}$ (1)
In Equation 1, the vectors $\vec{r}_{0}$ and $\vec{v}_{0}$ are the initial position and velocity vectors of the object. The universal variable $\chi$ is defined to be $\chi\equiv\sqrt{a}E$ for elliptical orbits and $\chi\equiv\sqrt{-a}F$ for hyperbolic orbits with semimajor axis $a$ and eccentric/hyperbolic anomaly $E$/ $F$. In Equation 1, $G$ is the gravitational constant, $M_{\odot}$ is the mass of the Sun, and $S(x)$ and $C(x)$ are Stumpff functions, defined as:
$\begin{split}S(x)&\equiv\begin{cases}\displaystyle\frac{\sqrt{x}-\sin\!\bigl{(}\sqrt{x}\bigr{)}}{\bigl{(}\sqrt{x}\bigr{)}^{3}},&x>0\ (\text{elliptical}),\\ \displaystyle\frac{1}{6},&x=0\ (\text{parabolic}),\\ \displaystyle\frac{\sinh\!\bigl{(}\sqrt{-x}\bigr{)}-\sqrt{-x}}{\bigl{(}\sqrt{-x}\bigr{)}^{3}},&x<0\ (\text{hyperbolic}),\end{cases}\end{split}$ (2)
and
$\begin{split}C(x)&=\begin{cases}\displaystyle\frac{1-\cos\!\bigl{(}\sqrt{x}\bigr{)}}{x},&x>0\quad(\text{elliptical}),\\ \displaystyle\frac{1}{2},&x=0\quad(\text{parabolic}),\\ \displaystyle\frac{\cosh\!\bigl{(}\sqrt{-x}\bigr{)}-1}{-\,x},&x<0\quad(\text{hyperbolic}).\end{cases}\end{split}$ (3)
For the application of the universal variable formalism to Kepler’s equations, Equations 2 and 3 are typically defined as functions of the variable $x=\chi^{2}/a$.
We calculate the numerical values of the universal variable $\chi$ by finding the roots of the transcendental equation defined by $\eta=0$. Following this step, we use the Lagrange propagators $f(\chi)$ and $g(\chi)$ (and their time derivatives) to advance the interceptor’s state to time $\tau_{0}+\Delta\tau$, since
$\vec{r}(\chi)=f(\chi)\,\vec{r}_{0}+g(\chi)\vec{v}_{0}\,,$ (4)
and
$\vec{v}(\chi)=\dot{f}(\chi)\,\vec{r}_{0}+\dot{g}(\chi)\vec{v}_{0}\,.$ (5)
We refer the reader to the corresponding Equations 25-28 in D. Seligman & G. Laughlin (2018) for definitions of the Lagrange propagators.
Since the arrival position $\vec{r_{T}}$ is taken directly from Horizons, we reformulate the boundary-value problem in terms of the variable $z=\chi^{2}/a$. With this simplification, there is no need to solve the target’s universal variable equation as described in D. Seligman & G. Laughlin (2018).
For a given interceptor state $\vec{r}_{0,I}$, $\vec{v}_{0,I}$, and $\Delta\tau$, the universal variable time-of-flight residual is given by:
$\begin{split}\eta_{I}\big{(}\chi_{I},\tau_{0}+\Delta\tau,\vec{r}_{0,I},\vec{v}_{0}+\vec{v}_{0,I}\big{)}={\lVert\vec{r}_{0,I}\rVert}\chi_{I}-\sqrt{GM_{\odot}}\Delta\tau\\ +\frac{{\vec{r}_{0,I}}\cdot(\vec{v}_{0}+\vec{v}_{0,I})}{\sqrt{GM_{\odot}}}\,\chi_{I}^{2}\,C\bigg{(}\frac{\chi_{I}^{2}}{a}\bigg{)}\\ +\left(1-\frac{\lVert\vec{r}_{0,I}\rVert}{a}\right)\chi_{I}^{3}S\bigg{(}\frac{\chi_{I}^{2}}{a}\bigg{)}\,.\end{split}$ (6)
Rather than propagating both orbits, we require that the interceptor reaches the known target position $\vec{r_{T}}$, (given by Horizons) after a flight time of $\Delta\tau$. To this end, we compute the angle between the target and interceptor, $\Delta\theta$, given by:
$\begin{split}\cos\big{(}\Delta\theta\big{)}&=\frac{\vec{r}_{T}\cdot\!\vec{r}_{I}}{\lVert\vec{r}_{T}\rVert\,\lVert\vec{r}_{I}\rVert}\,.\end{split}$ (7)
We define a geometric coefficient:
$\begin{split}A&=\sin\big{(}\Delta\theta\big{)}\,\sqrt{\frac{\lVert\vec{r}_{T}\rVert\,\lVert\vec{r}_{I}\rVert}{1-\cos\big{(}\Delta\theta\big{)}}}\,,\end{split}$ (8)
which rescales the chord connecting the departure and arrival position vectors so that the universal variable time-of-flight equation properly reflects the transfer geometry. We then introduce the intermediate function:
$\begin{split}y(z)&=\lVert\vec{r}_{T}\rVert+\lVert\vec{r}_{I}\rVert+A\,\frac{z\,S(z)-1}{\sqrt{C(z)}}\,.\end{split}$ (9)
With Equation 9, the time-of-flight residual can be written as defined in (H. D. Curtis, 2020, Chapter 5, Eq. 5.40):
$\begin{split}F(z)&=\Bigl{(}\frac{y(z)}{C(z)}\Bigr{)}^{3/2}\!S(z)+A\,\sqrt{y(z)}-\sqrt{{GM_{\odot}}}\,\Delta\tau\,,\end{split}$ (10)
Root-finding on Equation 10 enforces both geometric and temporal constrains simultaneously. For a root $z$, the Lambert-Lagrange coefficients are:
$\begin{split}f_{I}&=1-\frac{y(z)}{\lVert\vec{r}_{0,I}\rVert},\\ g_{I}&=A\,\frac{\sqrt{y(z)}}{\sqrt{GM_{\odot}}}.\end{split}$ (11)
By substituting these values into Equations (4, 5), the required arrival position and velocity vectors can be written as:
$\vec{r}_{T}=f_{I}\vec{r}_{0,I}+g_{I}\vec{v}_{0,I},$ (12)
and
$\begin{split}\vec{v}_{0,I}&=\frac{\vec{r}_{T}-f_{I}\vec{r}_{0,I}}{g_{I}}\,.\end{split}$ (13)
To evaluate the impulsive change in velocity $\Delta V$ required for a spacecraft to reach 3I/ATLAS, we subtract the heliocentric velocity of the departure planet from $\vec{v}_{0,I}$, so that:
$\begin{split}\Delta V&=\lVert\vec{v}_{0,I}-\vec{v}_{h}\bigr{\rVert}\,.\end{split}$ (14)
In Equation 14, $\vec{v}_{h}$ is the heliocentric velocity of departure planet.
We repeat the following process iteratively to solve for $\Delta V$:
  1. 1.
    Choose a trial departure epoch $\tau_{0,I}$ and set a flight time $\Delta\tau$.
  2. 2.
    Solve the universal variable Lambert problem for that $\Delta\tau$ by finding the root $z$ of the time-of-flight residual defined in Equation 10.
  3. 3.
    Compute the Lagrange coefficient using Equation 11 and propagate the interceptor state with Equations 12 and 13.
  4. 4.
    Evaluate $\Delta V$ using Equation 14.
  5. 5.
    Identify the global minimum-energy trajectory from the resulting $\Delta V$ values.
It is important to note that these $\Delta V$ values only encapsulate the impulsive burn required to leave a circular low-Earth orbit on an escape trajectory. In other words, they do not include the velocity required to leave the gravitational potential of the departure planet.
Finally, we verified that this methodology accurately produced flyby trajectories. We performed numerical integrations with hypothetical massless test particles representing the spacecraft and the Sun-Earth system with the solution velocity vectors. We verified that the position of the spacecraft and target matched after the corresponding time of flight. All verification numerical simulations were performed with REBOUND N-body code (H. Rein & S. F. Liu, 2012), the REBOUNDx library (D. Tamayo et al., 2020) and MERCURIUS (H. Rein et al., 2019).
Figure 1: Direct flight $\Delta V$ values as a function of launch date for trajectories to 3I/ATLAS from Earth. The color of each point corresponds to the flight time of the mission. The upper- and lower-panels show direct flight $\Delta V$ values from 2025 January 1 – 2026 March 31 and May 1 – November 2 2025, respectively. The minimum post-discovery $\Delta V$ trajectory is 24.0 km s-1 on 2025 July 1 with a flight time of 137 days. Pre-discovery trajectories may have been feasible, requiring $\Delta V\sim 7$ km s-1 with a flight times of $\sim$250 days. The shaded region on the upper panel is the region shown in the lower panel.
3 Optimal Trajectories from Earth
In this section, we examine the feasibility of trajectories from the Earth to the interstellar object 3I/ATLAS. We calculate minimum-energy trajectories from Earth to 3I/ATLAS using the methodology described in Section 2.
To determine when a direct launch from Earth could be viable, we use the Earth’s position as the initial launch date for a spacecraft. We then calculate the minimum $\Delta V$ required to reach 3I/ATLAS for a range of possible initial launch dates. The minimum $\Delta V$ required as a function of launch date for a direct flight is shown in Figure 1. The color of each point indicates the flight time for each of the corresponding trajectories.
The minimum $\Delta V$ solution recovered after the discovery of the object technically occurs on 2025 July 1. In other words, the optimal intercept mission would have required immediate launch of a spacecraft contemporaneous with discovery. This optimal trajectory would have required $\Delta V\sim 24.0$ km s-1, with a flight time of 137 days. The details of this trajectory are reported in Table 1. An orbit schematic of this hypothetical minimum-energy flyby mission is shown in Figure 2.
Table 1: Velocity vector information of minimum-energy trajectories to 3I/ATLAS for a variety of launch dates and locations. All listed velocities are with respect to the reference frame comoving with the orbit of the Earth.
Launch Date Launch Location $V_{x}$ $V_{y}$ $V_{z}$ $\Delta V$ Flight time Arrival Date
[km s-1] [km s-1] [km s-1] [km s-1] [days]
2025 January 10 Earth -5.353 0.592 4.371 6.935 248 2025 September 15
2025 July 1 Earth -8.196 22.536 -0.996 24.001 137 2025 November 15
2025 December 15 Earth 8.424 70.532 -4.082 71.151 198 2026 July 1
2025 March 6 Mars -0.483 -0.082 1.958 2.019 212 2025 October 4
2025 July 1 Mars -1.641 -1.033 2.959 3.538 94 2025 October 3
2025 August 10 Mars -3.291 -1.994 4.8354 6.179 54 2025 October 3
2025 November 10 Mars 14.410 72.571 3.959 74.093 233 2026 July 1
Figure 2: The orbit of a post-discovery minimum $\Delta V$ mission to 3I/ATLAS sent on 2025 July 1 from the Earth. The spacecraft would encounter 3I/ATLAS on 2025 November 15 and would require $\Delta V=24$ km s-1. The trajectory for 3I/ATLAS, Earth, and the spacecraft are plotted in purple, blue, and gray, respectively. Markers represent departure location and flyby location, and arrows show direction of each orbit/trajectory. Points on each trajectory indicate the positions of each object every 30 days.
The energetic requirements for a direct transfer mission post-discovery gradually increase during the timespan between July and early October 2025. Specifically the $\Delta V$ value increases to $\sim 35$ km s-1 while the flight time extends to $\sim 140$ days. After early October 2025 — but still before perihelion — the minimum $\Delta V$ rises steeply. The minimum $\Delta V$ reaches a value of $\sim 60$ km s-1 by December 2025 and $\sim 80$ km s-1 in January 2026. Therefore, an Earth-based direct flight is significantly more attainable if a spacecraft is launched before 3I/ATLAS passes its perihelion. However, even $\Delta V=24$ km s-1 for a fiducial best case mission shown in Figure 2 would be challenging for current propulsion capabilities.
A direct-transfer trajectory from Earth would have been more energetically feasible if 3I/ATLAS had been discovered earlier. Earth departures between January and June would have required $\Delta V\sim 10$ km s-1, with a flight time of approximately 150–200 days. A launch on 2025 January 10 would require only a minimum $\Delta V$ of 6.94 km s-1 with a flight time of 248 days, which may have been within the reach of modern spacecraft. It is worth noting that in a recent and complimentary analyses, A. Hibberd et al. (2025) identified an optimal trajectory from Earth to 3I/ATLAS. They calculated a minimum energy trajectory launched in 2024 July 19 for a rendezvous 2025 October 18. However, that trajectory assumed a continuous acceleration as opposed to a single impulsive burn.
Figure 3: Similar to Figure 1, but for trajectories to 3I/ATLAS from Mars. Here, the shaded inset region is from 2025 July 1 to 2025 September 25. The minimum energy trajectory for a post-discovery mission would be launched on 2025 July 1 with a flight time of 94 days and require $\Delta V=$ 3.54 km s-1. Pre-discovery trajectories would have required $\Delta V\sim 3$ km s-1 with flight times of more than 200 days.
Figure 4: Similar to Figure 2, except for a minimum $\Delta$V mission to 3I/ATLAS post-discovery, sent on 2025 July 1 from Mars. This hypothetical trajectory requires a $\Delta V\sim 4$ km s-1 and a time of flight of 94 days.
4 Optimal Trajectories from Mars
Mars has been a high-priority planetary science target for multiple space agencies, including NASA and the ESA, for several decades. A typical mission starts with the “prime mission”, which is designed to achieve the originally-funded objectives. Once a prime mission is completed, if the spacecraft can continue operating for a longer period of time, extend missions can be approved with additional science goals. As a spacecraft approaches the end of its operational life (e.g., running out of fuel) an end-of-mission maneuver is designed to safely dispose of the spacecraft. This end-of-mission maneuver can also be used to achieve additional objectives beyond the spacecraft’s initial design capabilities. For example, the Cassini spacecraft was disposed of by sending the spacecraft into Saturn’s atmosphere, taking in situ measurements that could not be accomplished during the active mission and which resulted in major scientific achievements (S. G. Edgington & L. J. Spilker, 2016). Additionally, the Rosetta mission ended its life by descending to the surface of its target comet 67P/Churyumov-Gerasimenko and making observations directly at the surface (A. Accomazzo et al., 2017; M. A. Barucci & M. Fulchignoni, 2017).
At this time, several spacecraft currently in orbit around Mars are approaching their end-of-life (M. H. Carr, 1996). If the required $\Delta V$ is small, one of these spacecraft could de-orbit from Mars and conduct a flyby of 3I/ATLAS. Such a maneuver would be a potentially high impact end-of-mission scenario for an end-of-life spacecraft. In this section, we assess whether trajectory designs from Mars to 3I/ATLAS may be energetically feasible. Although we consider generic trajectories, our results can be applied to existing spacecraft. We once again use the algorithm discussed in Section 2 to calculate the minimum-energy trajectories between 3I/ATLAS and Mars.
In Figure 3, we show the minimum $\Delta V$ required for such a mission. The minimum $\Delta V$ of 2.02 km s-1 occurs for a launch on 2025 March 6 (four months prior to discovery), with flight time of 212 days. The first post-discovery opportunity also occurs contemporaneously with discovery and would have required a $\Delta V$ of 3.54 km s-1 on 2025 July 1. A mission with this trajectory would arrive on 2025 October 3, three months after discovery, with its 94 days in transit. This optimal mission would fly by 3I/ATLAS during its closest approach to Mars. In Figure 4, we show the orbit of this minimum-energy post-discovery mission. For detailed information regarding these trajectories, see Table 1. After discovery, the required $\Delta V$ for flyby climbs gradually, yet remains at $\Delta V\leq 10$ km s -1 through 2025 August 30.
The sharp rise and gap in $\Delta V$ around October 2025 coincides with the closest approach between Mars and 3I/ATLAS, which occurs on 2025 October 3. The optimum trajectories sent prior to the closest approach have flight times such that the flyby occurs close in time to 2025 October 3. In Figure 5, we show the trajectory for a launch on 2025 August 10, 41 days after discovery, which demonstrates one of possible Mars departure scenarios. On the other hand, even the optimal trajectories after the close approach date would require significantly more $\Delta V$ to perform a flyby.
These results suggest that an intercept or flyby from Mars would require less $\Delta V$ than an Earth-based mission. A hypothetical Mars-based mission to 3I/ATLAS would have $\Delta V$ requirements within the performance envelope of modern spacecraft propulsion systems. Therefore, such a Mars-based flyby mission could be feasible.
Figure 5: Similar to Figure 2, except for a minimum-energy mission to 3I/ATLAS sent from Mars on 2025 August 10. Such a mission would require a $\Delta V\sim 6.5$ km s-1 and performs a flyby on 2025 October 4 with a time of flight of 54 days.
5 Discussion
In this paper, we calculated minimum-energy trajectories from both Earth and Mars to the interstellar object 3I/ATLAS. Given the current state of spacecraft propulsion technology, a spacecraft flyby of 3I/ATLAS would be more feasible if launched from Mars rather than Earth.
Our calculations show that post-discovery Earth departures likely demand $\Delta V$ values beyond the capabilities of available spacecraft. However, had 3I/ATLAS been discovered before July 1, a mission launched from Earth would have required $\Delta V\sim 7$ km s-1, within the envelope of current spacecraft technologies. This fact underscores the critical impact of discovery time on the feasibility of deep-space missions.
On the other hand, Mars departures require substantially less $\Delta V$. For example, a pre-discovery flyby on 2025 March 6 (four months prior to discovery) could reach 3I/ATLAS with just $\Delta V=2.02$ km s-1 while a launch on the discovery date of 2025 July 1 would require $\Delta V\sim 4$ km s-1. These $\Delta V$ values likely fall within the maneuvering envelope of current Mars orbiters, making it possible to re-target one for a close flyby of 3I/ATLAS.
There are a variety of spacecraft currently orbiting Mars that are nearing their end of life — Mars Reconnaissance Orbiter (R. W. Zurek & S. E. Smrekar, 2007), MAVEN (B. M. Jakosky et al., 2015), Mars Odyssey (R. S. Saunders et al., 2004), Mars Express (A. Chicarro et al., 2004; A. F. Chicarro, 2007), ExoMars Trace Gas Orbiter (A. C. Vandaele et al., 2015) and others. These spacecraft collectively carry a wide range of imaging cameras, spectrometers, in situ instrumentation, and communications and operation tools. Repurposing these missions as interstellar interceptors would offer valuable scientific data, especially since they may be able to reach interstellar objects that are inaccessible from Earth. While repurposing Earth-observing missions is also possible, these spacecraft are generally de-orbited during their end-of-life phase and have less maneuvering capability, limiting their use as interstellar interceptors.
The Janus spacecraft mission (D. J. Scheeres et al., 2021), part of NASA’s SIMPLEx program, consists of two identical 12 U CubeSats, designed to fly by binary asteroids. Each carries JCam, a visible/infrared imager capable of capturing meter-scale resolution images and thermal data. Janus could have potentially performed its 10 km s-1 flyby of 3I/ATLAS if we had launched it within the first few days after discovery since it is one of the ready-to-launch spacecraft. This close encounter would allow Janus to collect detailed images and thermal data, helping us to understand 3I/ATLAS’s shape, rotation, and dust activity.
Although a flyby may not be feasible, NASA’s OSIRIS-APEX (D. N. DellaGiustina et al., 2023) could be assigned to observe 3I/ATLAS from the spacecraft’s existing trajectory using the PolyCam and MapCam imagers (B. Rizk et al., 2018). Observations with the Eight Color Asteroid Survey (ECAS) $b$, $v$, $w$, and $x$ filters onboard the OSIRIS-APEX MapCam instrument (D. N. DellaGiustina et al., 2018) would provide valuable data characterizing 3I/ATLAS. MapCam could acquire narrow-band photometry from the near-ultraviolet through the near-infrared, constraining the spectral slope and broad surface composition of 3I/ATLAS (D. R. Golish et al., 2021; D. N. DellaGiustina et al., 2023). The viable observation windows are late September and mid-November of 2025. Although late September offers the advantage of pre-perihelion observations, preferable for comet characterization, it coincides with an OSIRIS-APEX Earth Gravity Assist (D. N. DellaGiustina et al., 2023), making additional spacecraft activities during this period potentially risky. By mid-November, the solar elongation of 3I/ATLAS increases rapidly, and observations could likely be conducted once the elongation exceeds 35–40 degrees. A spacecraft like OSIRIS-APEX can provide valuable photometric data at a distance, obtaining unique observation geometries in the absence of atmospheric effects.
In addition, higher-resolution imagery could resolve any dust plumes or jets coming off the surface of 3I/ATLAS, further characterizing its mass-loss rate. By viewing the object at different phase angles, the albedo of the object can be constrained (assuming that we can resolve the nucleus). Finally, a simple infrared measurement would indicate how the surface of 3I/ATLAS responds to thermal forcing, enabling further characterization of its surface properties. Together, these results would allow for estimates of nucleus size, composition, rotation, and activity level.
Even if no existing spacecraft performs a flyby of 3I/ATLAS, the discovery and apparition provide a tangible example for a future interstellar object mission. The NSF-DOE Vera C. Rubin Observatory Legacy Survey of Space and Time (LSST) should identify more interstellar objects in the near future (A. Moro-Martín et al., 2009; N. V. Cook et al., 2016; T. Engelhardt et al., 2017; D. J. Hoover et al., 2022; D. Marčeta, 2023; R. C. Dorsey et al., 2025). If we station fueled spacecraft at key locations (e.g., Earth orbit, Mars orbit, the Earth-Sun Lagrange points) we could direct these craft to fly by a newly discovered interstellar object with short notice. If equipped with narrow‑band cameras and sufficient maneuvering capability for small trajectory changes, these fueled spacecraft could execute fast flybys or continuous brightness monitoring of interstellar objects immediately after discovery. Such measurements would provide critical information regarding the composition, rotational state, and activity of interstellar objects beyond the capabilities of ground-based instruments alone. The data provided by such a mission would provide unprecedented insights into the interstellar object population, and in turn, the history of planet formation throughout the Galaxy.
6 acknowledgments
We thank Shannon Curry, Adina Feinstein, Karen Meech, James Wray, Avi Loeb, Devin Hoover, and Qicheng Zhang for helpful conversations.
D.Z.S. is supported by an NSF Astronomy and Astrophysics Postdoctoral Fellowship under award AST-2303553. This research award is partially funded by a generous gift of Charles Simonyi to the NSF Division of Astronomical Sciences. The award is made in recognition of significant contributions to Rubin Observatory’s Legacy Survey of Space and Time. A.Y. and T.F. also acknowledge support from NSF grant number AST-2303553. A.G.T. acknowledges support from the Fannie and John Hertz Foundation and the University of Michigan’s Rackham Merit Fellowship Program. E.P.-A. acknowledges support by the Italian Space Agency (ASI) within the LUMIO project (ASI-PoliMi agreement n. 2024-6-HH.0). K.E.M. acknowledges support from the NASA ROSES Discovery Data Analysis Program (DDAP). D.N.D. and M.C.N. acknowledge support under NASA contract NNM10AA11C.